OpenStudy (anonymous):

need help with this one question x 2/3 divided by x 1/5 the answer needs to be in radical form...please show me the steps on how you got your answer

OpenStudy (anonymous):

hello i really need sombody help on this question

OpenStudy (anonymous):

could you help me out here @amoodarya

OpenStudy (anonymous):

@dpaInc

OpenStudy (anonymous):

@Ashleyisakitty

OpenStudy (amoodarya):

do you mean 2 x* -- 3 ----- = ? 1 x*--- 5

OpenStudy (anonymous):

yes

OpenStudy (amoodarya):

\[x ^{\frac{ 2 }{ 3 }}\div x ^{\frac{1 }{ 5 }}=x^{\frac{ 2 }{ 3 }-\frac{1 }{ 5 }}\] or this one ?

OpenStudy (anonymous):

its the first one

OpenStudy (anonymous):

its a division problem

OpenStudy (amoodarya):

\[\frac{2 }{ 3 }-\frac{1 }{ 5 }=\frac{2\times 5-1\times3 }{ 3\times5 }=\frac{7 }{ 15 }\] \[x^{\frac{7 }{15 }}=\sqrt[15]{x^7}\]

OpenStudy (amoodarya):

for the first we have simplify x so 2 -- 3 10 ---- = --- 1 3 -- 5

OpenStudy (anonymous):

where did you get the 10/3 from i thought you was suppose to find the gcf

OpenStudy (amoodarya):

far by far , near by near 2 -- 3 2*5 ----=----- 1 3*1 -- 5

OpenStudy (anonymous):

oh ok i got you

OpenStudy (anonymous):

and the problem you had did in the beginning is that the answer?

OpenStudy (amoodarya):

yes , if your problem was the first

OpenStudy (anonymous):

it was a subtraction problem it was a division prob

OpenStudy (anonymous):

@amoodarya